OpenStudy (anonymous):

The value of -7 + 24 ÷ (-3)(2) is _____.

OpenStudy (gorv):

first we will open the bracket

OpenStudy (anonymous):

ok i did

OpenStudy (anonymous):

whats the answer thoe

OpenStudy (anonymous):

-23

OpenStudy (gorv):

-7+24/(-3)(2)=-7+24/(-6)=-7-4=-11

OpenStudy (gorv):

if bracket wasnt there in between than it would have been -23